OpenStudy (anonymous):

IS the derivative of y=x^2*Cot(x)-1/x^2 is equal to 2x*cot(x)-x^2*csc(x)-2/x^3?

OpenStudy (anonymous):

i got this answer by my self, and my ti-89 got \[\frac{ 2x^4\sin(X)Cos(x)+2(\sin(x))^2-x^5 }{x^3(\sin(x))^2 }\]

OpenStudy (psymon):

\[x^{2}\cot(x) -\frac{ 1 }{ x^{2} }\] correct?

OpenStudy (anonymous):

yes sir!

OpenStudy (psymon):

Alrighty, cool. So let's see, I get: 2xcot(x) - x^2csc^2(x)+ 2/(x^3) Should be something like that O.o

OpenStudy (anonymous):

doing it by my self i got \[2xcot(x)-x^2cscx-\frac{ 2 }{ x^3 }\]

OpenStudy (psymon):

So same thing I get but wrong sign on the last part xD

OpenStudy (anonymous):

yeah i got the same thing, but my ti is throwing that answer to me

OpenStudy (anonymous):

why you got it possitve?

OpenStudy (anonymous):

ooo it was already negative so negative times negative equal positive

OpenStudy (psymon):

Well, the ti LOVES to turn things into sin and cos, so that makes it confusing. And if I had to go through it the long way: \[-\frac{ 1 }{ x^{2} }=-x^{-2} \implies -(-2)x^{-3} = \frac{ 2 }{ x^{3} }\]

OpenStudy (anonymous):

cool thanks!!

OpenStudy (psymon):

Yeah, sure :3

OpenStudy (anonymous):

so the answer the TI gave me is correct just in the long way?

OpenStudy (psymon):

Likely. In fact, it looks like it turned cot into cos/sin and did quotient rule, then combined everything into one common denominator, lol. It just made a mess of things.

OpenStudy (anonymous):

damn

OpenStudy (psymon):

Yeah, thatll be the bad thing about checking derivatives with it x_x

OpenStudy (anonymous):

hey what is the derivative of (sin(x))*(Tan(x)). cosx*sec^2x?

OpenStudy (psymon):

sin times tan? cos(x)tan(x) + sin(x)sec^2(x)

OpenStudy (anonymous):

derivative of (sin(x))*(Tan(x)). My answer is cosx*sec^2x

OpenStudy (anonymous):

yeah sin times tan

OpenStudy (psymon):

Yeah, should be right what I put.

OpenStudy (anonymous):

yup your right!

OpenStudy (psymon):

I sped through it, so gotta double check xD

OpenStudy (anonymous):

i check it and i say is correct.

OpenStudy (anonymous):

and like always my fluttering TI giving other pellet

OpenStudy (anonymous):

can always use: http://www.wolframalpha.com/ to check derivative answers too

OpenStudy (anonymous):

are you there psymon?

OpenStudy (psymon):

Yep.

OpenStudy (anonymous):

i have a new one, \[G(x)= Csc(x)Cot(x)\]

OpenStudy (psymon):

Too bad thats not integral of, then its easy, lol.

OpenStudy (anonymous):

and my answer is \[\cot(x)(-\csc(x)\cot(x))-\csc(x)(\csc^2x)\]

OpenStudy (anonymous):

should i do something? like add the cot(x)'s and the csc(x)'s?

OpenStudy (psymon):

-cot(x)csc(x)cot(x) - csc(x)csc^2(x) So yeah, looks the same, just combine what you can from there.

OpenStudy (anonymous):

Yei, this isn't hard at all my TI and my professor makes it complicated. damn

OpenStudy (psymon):

Lol, yeah, if you are used to doing them, theyre really simple. But just like everything else when you get to calculus level, its more tedious and possible to mess up than it is hard.

OpenStudy (anonymous):

i got other one \[y= \frac{ \cot x }{1+ \cot x }\]. let me get the answer

OpenStudy (psymon):

This is one id so cheat on, haha.

OpenStudy (anonymous):

the answer is \[-\frac{ Csc^2 x }{ 1-\csc^2x }\]

OpenStudy (psymon):

Thats your answer, right?

OpenStudy (anonymous):

i think is 1

OpenStudy (anonymous):

yeah i got that

OpenStudy (psymon):

Possible our answers are the same, but I do that kind of derivative in a sneaky way.

OpenStudy (anonymous):

how you do it then?

OpenStudy (psymon):

Well, doing it the normal way, I get: \[-\frac{ \csc^{2}x }{ (1+cotx)^{2} }\]

OpenStudy (anonymous):

explain me why numerator got possitive and to the second power, pleaasseee:D

OpenStudy (psymon):

Yeah. Well, if you do quotient rule, the denomintor is supposed to be squared normally. I assume you know the quotient rule, so Ill just write out what you have after the derivative of everything. \[\frac{ -\csc^{2}x(1+cotx) + \cot(x)\csc^{2}x }{ (1+cotx)^{2} }\]

OpenStudy (anonymous):

oooo i got it

OpenStudy (psymon):

Yep. But I can show you how I do it a tricky way if you want.

OpenStudy (anonymous):

yeah that would be could! please

OpenStudy (psymon):

\[\frac{ cotx }{ 1+cotx }\] Add 1 and subtract 1 at the same time in the numerator. This is a trick where you technically add 0, but it allows you to make more than one fraction. So I do this: \[\frac{ cotx + 1 - 1 }{ 1+cotx }\] Now you are allowed to make a new fraction for every term in the numerator of a fraction or any combination of terms in the numerator. So that means I can rewrite the above like this: \[\frac{ cotx + 1 }{ 1+cotx }- \frac{ 1 }{ 1+cotx }\implies 1 - \frac{ 1 }{ 1+cotx }\] So now if I take the derivative, the 1 goes away completely and the other fraction becomes \[-(1+cotx)^{-1} \implies (1+cotx)^{-2}*-\csc^{2}x \implies \frac{ -\csc^{2}x }{ (1+cotx)^{2} }\]

OpenStudy (anonymous):

heyyyy! good trick ! let me try it in other

OpenStudy (psymon):

Lol, alright then xD

OpenStudy (anonymous):

this time is \[y=\frac{ \cos x}{ 1+\sin x }\]

OpenStudy (anonymous):

damn i got stuck in \[\frac{ \cos x +1 }{ 1 +\sin x } - \frac{1 }{ 1 +\sin x }\]

OpenStudy (psymon):

Yeah, that would work. Its just having cosx+1/1+sinx doesnt do you much good, lol. Probably not a situation where you want to do + 1 - 1

OpenStudy (anonymous):

should i use quotient rule?

OpenStudy (psymon):

Quotient or product rule, you can do either one. But yeah, no real super trick for this one xD

OpenStudy (anonymous):

hey i got \[-\frac{ \cos x \cos x }{(1+\sin x) ^2 }\]. is it correct?

OpenStudy (psymon):

I dont get that, no.

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

what about\[\frac{ (- \cos x)^2+\sin (x)(\sin (x) +1) }{ (1+\sin (x))^2 }\]

OpenStudy (psymon):

I did product rule, so let me check that method.

OpenStudy (psymon):

Looks like youre missing a negative. Gah, hating doing it the quotient rule way xD \[\frac{ -sinx(1+sinx) - \cos(x)\cos(x) }{ (1+sinx)^{2} }\] \[\frac{ -sinx-\sin^{2}x-\cos^{2}x }{ (1+sinx)^{2} }\] \[\frac{ -sinx-(\sin^{2}x+\cos^{2}x) }{ (1+sinx)^{2} }\] \[\frac{ -sinx-1 }{ (1+sinx)^{2} }=\frac{ -(sinx+1) }{ (sinx+1)^{2} }\implies \frac{ -1 }{ (sinx+1) }\]

OpenStudy (anonymous):

where the sin^2 x come frome, in the second fraction

OpenStudy (psymon):

I multiplied the -sinx through. -sinx(1+sinx) = -sinx-sin^2(x)

OpenStudy (anonymous):

got it!

OpenStudy (psymon):

Lol, alright, cool xD

OpenStudy (anonymous):

\[\frac{−sinx−(\sin2x+\cos2x) }{ (1+sinx)2 }\] so this suppose to be the answer?

OpenStudy (anonymous):

or \[\frac{ −1−(\sin^2x+\cos^2x) }{ (1+sinx)2 }\]

OpenStudy (psymon):

You can keep going like I did. Can you follow what I did?

OpenStudy (anonymous):

Yes i do,

OpenStudy (psymon):

Okay, cool. Yeah, so the answer turns out to be something pretty simplified it seems.

OpenStudy (anonymous):

what an idiot of me, i wasnt doing the sum of sin^2x+ cos^x

OpenStudy (psymon):

Yeah, simplifies to 1, lol.

OpenStudy (anonymous):

what about \[ Y= X^2 \sin x + 2x \cos x - 2 \sin x\]

OpenStudy (anonymous):

i used the sum and difference rule, which it gave me the result of \[2x \cos x - 2 \sin x- cosx\]

OpenStudy (anonymous):

ups\[2xcosx−2sinx+cosx\]

OpenStudy (psymon):

Thats a new problem, right?

OpenStudy (anonymous):

the new problem is\[Y=X^2sinx+2xcosx−2sinx\]

OpenStudy (psymon):

2xsinx + x^2cosx + 2cosx -2xsinx - 2cosx Everything cancels to x^2cosx.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

i know the answer is x^2cosx but where did x^2cosx come from?

OpenStudy (anonymous):

oooooooo i get it, you differentiate each value by its own?

OpenStudy (psymon):

Yeah. Kinda have to. I mean, there may be things you can do, but I go term by term. First term is product rule, 2nd term is another product rule, last one is straightforward.

OpenStudy (anonymous):

straightfoward?

OpenStudy (psymon):

As in nothing tricky. It just is a simple derivative.

OpenStudy (anonymous):

what about this one \[Y=x^2Cosx-2xSinx-2cosx\]

OpenStudy (anonymous):

and my answer is\[-x^2sinx\]

OpenStudy (psymon):

Yeah, it even looks the same: 2xcosx - x^2sinx - 2sinx -2xcosx + 2sinx simplifies to -x^2sinx

OpenStudy (anonymous):

looks the same as mine

OpenStudy (psymon):

Yeah, guess so xD I get the same answer.

OpenStudy (anonymous):

i need to remember this rules all of them:(

OpenStudy (psymon):

Yeah, remembering product rule and quotient rule and them being able to do them correctly. Takes practice x_x

OpenStudy (anonymous):

i have my test in three weeks and i dont know very well Limit theme, some of them are easy just by substituting, the others are hard as hell

OpenStudy (psymon):

Well, these arent limits, lol. If youre worried about those why not bring some of those up?

OpenStudy (anonymous):

because right know im doing some homework about trigonometric Function differentiation

OpenStudy (anonymous):

i want to bring some up, but homework first then i study limits

OpenStudy (psymon):

Ah. Fair enough, lol.